OpenStudy (anonymous):

What value of x makes the equation true? –8.6 = –4.1 + x A. –12.7 B. –4.5 C. 4.5 D. 12.7

OpenStudy (anonymous):

@PizzaLover123 and @the_ocean_girl

OpenStudy (alexandervonhumboldt2):

add 4.1 to both sides: -8.6+4.1=-4.1+x+4.1 -8.6+4.1=x simplify and get the answer

OpenStudy (anonymous):

thx